In a race from pt. X to pt Y and back, Jack averages 30miles/hr to pt Y and 10 miles/hr back to pr X.Sandy averages 20 miles/hr in both directions. If Jack and Sandy start race at same tym, who’ll finish 1st ?
Answer by Alan3354(69443)

In a race from pt. X to pt Y and back, Jack averages 30miles/hr to pt Y and 10 miles/hr back to pr X.Sandy averages 20 miles/hr in both directions. If Jack and Sandy start race at same tym, who’ll finish 1st ?
------------------
d = distance
Jack's time = d/30 + d/10 = 4d/30 = 2d/15 = 0.1333... d
Sandy's time = d/20 + d/20 = d/10 = 0.1d
----------------
This is the basis of the Michelson-Morley experiment that disproved the existence of the "ether", I think in the early 1900's.